OpenStudy (anonymous):

what is the solution of this equation? -5+7b=8b-6

OpenStudy (anonymous):

b=1

OpenStudy (anonymous):

Get all the terms with a 'b' in it to one side of the equals sign, and all the constant terms tot he other side.

OpenStudy (aravindg):

1

OpenStudy (anonymous):

b=1 , sorry mistyping

OpenStudy (anonymous):

\[-5+7b=8b-6\] \[-5+6=8b-7b\] \[1=b\]

OpenStudy (anonymous):

-5+7b=8b-6 =7b-8b=5-6 =-b=-1 b=1